OpenStudy (anonymous):

The eqns and enthalpy changes for two reactions used in the manufacture of sulphuric acid are: S + O2 --> SO2 Delta H = -300 kJ 2SO2 + O2 --> 2SO3 Delta H = -200 kJ What is the enthalpy change, in kJ, for the reaction below? 2S + 3O2 --> 2SO3

OpenStudy (frostbite):

I see you did not understand it 100% anyway from yesterday, so I'm gonna go over this one very slowly: We are going to use Hess's law again. So we start we solve the problem like this: 1) we set up the reactions so they come on the form as the reaction we wish. 2) we indentify the single reactions enthalpy. 3) we use Hess's law to find the answer.

OpenStudy (chmvijay):

haa haaa haaa thats true said @Frostbite

OpenStudy (anonymous):

I understood yesterday's example but this seems different somehow! Thanks for coming to help!

OpenStudy (frostbite):

It isn't and you'll get to see it. _________ Part 1) The reaction we want: \[\Large 2 ~ S + 3 ~ O _{2} \to 2 ~ SO _{3}\] We start by writing up the first reaction: \[\Large S + O _{2} \to SO _{2}\] We multiply this reaction with 2 in the start: \[\Large 2 ~ S + 2 ~ O _{2} \to 2 SO _{2}\] We add the REVERSE reaction to this one: \[\Large 2 ~ S + 2 ~ O _{2} + 2 ~ SO _{2} + O _{2} \to 2 ~ SO _{3} + 2 ~ SO _{2}\] The \(SO_{2}\) cancels and \(2 ~ O_{2}+O_{2}=3 ~ O_{2} \). So we get: \[\Large 2 ~ S + 3 ~ O _{2} \to 2 ~ SO _{3}\] Just as we wanted. With me so far? :)

OpenStudy (frostbite):

Wait sorry I did not use the reverse reaction. my mistake I used it as it was.

OpenStudy (anonymous):

Ok so after you multiply eqn 1 by 2

OpenStudy (frostbite):

Then I add the second reaction. To "add" a reaction to another is when you combine the reactants for the reactions for them self, and the products for them self.

OpenStudy (anonymous):

Yay

OpenStudy (anonymous):

I get. It so, it goes -300 x 2 - 200

OpenStudy (anonymous):

So the answer is -800kJ

OpenStudy (frostbite):

exactly: \[\Large 2 ~ S + 2 ~ O _{2} + 2 ~ SO _{2} + O _{2} \to 2 ~ SO _{3} + 2 ~ SO _{2}\] \[\large \Delta H ^{\Theta}=2 \times \Delta H ^{\Theta}(1)+\Delta H ^{\Theta}(2)\] \[\large \Delta H ^{\Theta}=2 \times (-300 ~ kJ)+(-200 ~ kJ)=-800 ~ kJ\] Yea :)

OpenStudy (frostbite):

Not that hard anyway right? :)

OpenStudy (anonymous):

Nope, just gotta get used to it! Thanks again!

OpenStudy (chmvijay):

@amsterdam11 u got 100% or noo :) if u have understood u wont ask any more question on this i think :) bcoz u can solve them yourself :)

OpenStudy (anonymous):

I'm gonna try another one and if I can't do it, I'll be back but I think I got it!

OpenStudy (frostbite):

But well done @amsterdam11 glad you could finish the question before I was done explaining, show some good progress! :)

OpenStudy (anonymous):

Just did another q and got it right! Thank you to you lot!